# Exploration v.s. Exploitation: Rethinking RLVR through Clipping, Entropy, and Spurious Reward

### v1

@@ 4.1 One-step Policy Entropy Change under Random Rewards
${\mathbb{E}}[{\mathcal{H}}(\pi_{\textnormal{new}})-{\mathcal{H}}(\pi_{\textnormal{old}})]\leq-c_{G}\Phi(\pi_{\rm old})\eta^{2}+{\mathbb{E}}[R_{c}(\eta)]+c(p)G\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right),$
where $c_{G}$ and $\Phi$ are defined as the same as in Theorem˜4.1; for other constants, see Section˜C.3.
@@ C.3 Missing Proofs in §4
${\mathbb{E}}[{\mathcal{H}}(\pi_{\textnormal{new}})-{\mathcal{H}}(\pi_{\textnormal{old}})]=-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+\mathbb{E}[R(\eta)],\quad|R(\eta)|\leq C\eta^{4}.$
@@
with
@@
We rewrite the update rule as follows,
We define $\psi(\eta):=\log Z(\eta)$ and $u(a):=\eta\tilde{A}(a)-\psi(\eta)$. Then, we have $\pi_{\eta}(a)=\pi_{\textnormal{old}}(a)e^{u(a)}$ and
${\mathbb{E}}_{\pi_{\textnormal{old}}}[e^{u}]=\sum_{a}\pi_{\textnormal{old}}(a)e^{\eta\tilde{A}(a)-\psi(\eta)}=1.$
We write $W(a)=\log(\pi_{\textnormal{old}}(a))$. Then, the change of entropy satisfies
$\Delta{\mathcal{H}}(\eta)={\mathcal{H}}(\pi_{\eta})-{\mathcal{H}}(\pi_{\textnormal{old}})=-{\mathbb{E}}_{\pi_{\textnormal{old}}}[e^{u}(L+u)]+{\mathbb{E}}_{\pi_{\textnormal{old}}}[L].$
Let $f(\eta):={\mathbb{E}}[\Delta{\mathcal{H}}(\eta)]$, where $\mathbb{E}$$\cdot$] is taken over the randomness from $\{\mathbf{y}^{(i)}\}_{i=1}^{G}$ and $\{\mathbf{r}_{i}\}_{i=1}^{G}$.
Leading Term.
By Lemma˜2.4, $A_{i}$ is symmetric and thus $\tilde{A}(a)$ is symmetric. Thus, we have
@@
This implies that $f$ is even and $f^{\prime}(0)=f^{(3)}(0)=0$. One has the identity
$\tfrac{\mathrm{d}}{\mathrm{d}\eta}\left({\mathbb{E}}_{\pi_{\eta}}[g]\right)=\textnormal{Cov}_{\pi_{\eta}}(\tilde{A},g),\quad\textnormal{for any function }g.$
Using ${\mathcal{H}}(\pi_{\eta})=-{\mathbb{E}}_{\pi_{\eta}}[\log(\pi_{\eta})]$ and $\log(\pi_{\eta}(a))=L(a)+\eta\tilde{A}(a)-\psi(\eta)$, we have
[…]
and
${\mathcal{H}}^{\prime\prime}(\pi_{\eta})=-\textnormal{Cov}_{\pi_{\eta}}(\tilde{A}^{2},L)-\textnormal{Var}_{\pi_{\eta}}(\tilde{A})-\eta{\mathbb{E}}_{\pi_{\eta}}[(\tilde{A}-{\mathbb{E}}_{\pi_{\eta}}[\tilde{A}])^{3}],$
Evaluating at $\eta=0$ yields
[…]
Then, we have
[…]
@@
This implies
$f^{\prime\prime}(0)=-\tfrac{1-2^{1-G}}{G}\Phi(\pi_{\textnormal{old}}).$
Remainder Term.
Note that $f$ is even. For any $\eta>0$, there exists $t\in(0,\eta)$ such that
$f(\eta)=\tfrac{f^{\prime\prime}(0)}{2}\eta^{2}+\tfrac{f^{(4)}(t)}{24}\eta^{4}.$
It suffices to bound $\sup_{t\in[0,\eta]}|f^{(4)}(t)|$. Indeed, we let $X_{c}(a)=\tilde{A}(a)-{\mathbb{E}}_{\pi_{t}}[\tilde{A}$] and $L_{c}(a)=L(a)-\mathbb{E}_{\pi_{t}}[L]$. Then, we have
$\tfrac{\mathrm{d}^{4}}{\mathrm{d}t^{4}}{\mathcal{H}}(\pi_{t})=-{\mathbb{E}}_{\pi_{t}}[L_{c}X_{c}^{4}]+4{\mathbb{E}}_{\pi_{t}}[X_{c}^{3}]{\mathbb{E}}_{\pi_{t}}[L_{c}X_{c}]$
$+6{\mathbb{E}}_{\pi_{t}}[X_{c}^{2}]{\mathbb{E}}_{\pi_{t}}[LX_{c}^{2}]-3{\mathbb{E}}_{\pi_{t}}[X_{c}^{4}]+9({\mathbb{E}}_{\pi_{t}}[X_{c}^{2}])^{2}+t(10{\mathbb{E}}_{\pi_{t}}[X_{c}^{2}]{\mathbb{E}}_{\pi_{t}}[X_{c}^{3}]-{\mathbb{E}}_{\pi_{t}}[X_{c}^{5}]).$
We bound the right-hand side using ${\mathbb{E}}_{\pi_{t}}[X^{4}]$. Indeed, we have $L(a)=\log(\pi_{\textnormal{old}}(a))\in[\log(\pi_{\min}),0]$, so $\|L-{\mathbb{E}}_{\pi_{t}}[L]\|_{\infty}\leq\log(\frac{1}{\pi_{\min}})$. Thus, we have
$|-{\mathbb{E}}_{\pi_{t}}[L_{c}X_{c}^{4}]+4{\mathbb{E}}_{\pi_{t}}[X_{c}^{3}]{\mathbb{E}}_{\pi_{t}}[L_{c}X_{c}]+6{\mathbb{E}}_{\pi_{t}}[X_{c}^{2}]{\mathbb{E}}_{\pi_{t}}[LX_{c}^{2}]|\leq 11\log(\tfrac{1}{\pi_{\min}}){\mathbb{E}}_{\pi_{t}}[X_{c}^{4}].$
We also have
$\left|-3{\mathbb{E}}_{\pi_{t}}[X_{c}^{4}]+9({\mathbb{E}}_{\pi_{t}}[X_{c}^{2}])^{2}\right|\leq 12{\mathbb{E}}_{\pi_{t}}[X_{c}^{4}].$
Using ${\mathbb{E}}_{\pi_{t}}[|X_{c}|]\leq 2{\mathbb{E}}_{\pi_{t}}[|\tilde{A}(a)|]\leq\frac{1}{\pi_{\min}}$ (see the proof of Lemma˜2.4), we have
$\left|10{\mathbb{E}}_{\pi_{t}}[X_{c}^{2}]{\mathbb{E}}_{\pi_{t}}[X_{c}^{3}]-{\mathbb{E}}_{\pi_{t}}[X_{c}^{5}]\right|\leq\tfrac{11}{\pi_{\min}}{\mathbb{E}}_{\pi_{t}}[X_{c}^{4}].$
By using ${\mathbb{E}}_{\pi_{t}}[X_{c}^{4}]\leq 16{\mathbb{E}}_{\pi_{t}}[\tilde{A}^{4}]$, we conclude that for all $t\in[0,\eta]$,
$\left|\tfrac{\mathrm{d}^{4}}{\mathrm{d}t^{4}}{\mathcal{H}}(\pi_{t})\right|\leq\left(12+11\log(\tfrac{1}{\pi_{\min}})+\tfrac{11t}{\pi_{\min}}\right){\mathbb{E}}_{\pi_{t}}[X_{c}^{4}]\leq\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176t}{\pi_{\min}}\right)\mathbb{E}_{\pi_{t}}[\tilde{A}^{4}].$
In addition, we have
${\mathbb{E}}_{\pi_{t}}[\tilde{A}^{4}]\leq e^{t/\pi_{\min}}{\mathbb{E}}_{\pi_{\textnormal{old}}}[\tilde{A}^{4}].$
Putting these pieces together yields
$\sup_{t\in[0,\eta]}|f^{(4)}(t)|\leq e^{\eta/\pi_{\min}}\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176\eta}{\pi_{\min}}\right){\mathbb{E}}\left[{\mathbb{E}}_{\pi_{\textnormal{old}}}[\tilde{A}^{4}]\right].$
Conclusion.
We define $I_{i}(a)=\bm{1}\{\mathbf{y}^{(i)}=a\}$. Thus, we have $I_{i}(a)\sim\textnormal{Bernoulli}(p)$ i.i.d. and it is independent of $\{A_{i}\}_{i=1}^{G}$. We write $S(a)=\sum_{i=1}^{G}A_{i}I_{i}(a)$ and
@@ Conclusion.
$S_{1}$ $=\tfrac{|{\mathcal{V}}|-1}{\pi_{\min}}+\tfrac{1}{1-(|{\mathcal{V}}|-1)\pi_{\min}},$
$S_{2}$ $=\tfrac{|{\mathcal{V}}|-1}{(\pi_{\min})^{2}}+\tfrac{1}{(1-(|{\mathcal{V}}|-1)\pi_{\min})^{2}}.$
This completes the proof. ∎
@@
Finally, we show that $r_{c}(a)\leq 1+\varepsilon$ for all $a\in{\mathcal{V}}$ given Eq.˜17. Suppose that there are some $a^{\star}$ so that $r_{c}(a^{\star})>1+\varepsilon$. Then, by Lemma˜C.1, we have $\xi_{a^{\star}}=0$ and
@@
Putting these pieces together yields a direct contradiction to Eq.˜17 as follows,
@@
Lemma C.3.
[…]
$|{\mathbb{E}}[R(\eta)]|\leq(1-q_{\hat{\pi}})C(\hat{\pi})\eta^{4}+q_{\hat{\pi}}C(\pi_{\min})\eta^{4},$
where $C(\pi)$ is defined as the same as in Theorem˜4.1.
Proof.
We mimic the proof of Theorem˜4.1 (and using the same notation) to obtain:
$f(\eta)=\tfrac{f^{\prime\prime}(0)}{2}\eta^{2}+\tfrac{f^{(4)}(t)}{24}\eta^{4}.$
We reuse the expression derived for $\frac{\mathrm{d}^{4}}{\mathrm{d}t^{4}}{\mathcal{H}}(\pi_{t})$. Now, we split the expectation as follows,
${\mathbb{E}}[R(\eta)]={\mathbb{E}}[R(\eta)\mid B_{\hat{\pi}}^{c}](1-q_{\hat{\pi}})+{\mathbb{E}}[R(\eta)\mid B_{\hat{\pi}}]q_{\hat{\pi}}$.
On $B_{\hat{\pi}}^{c}$, we have $\pi_{\textnormal{old}}(a)\geq\hat{\pi}$ for all $a\in{\mathcal{S}}$. Since $\tilde{A}(a)=0$ outside $\mathcal{S}$}, every place in the remainder proof where $\pi_{\min}$ is used to upper bound denominators (e.g. $|\tilde{A}|,{\mathbb{E}}_{\pi_{t}}[\tilde{A}^{4}]$, and the change-of-measure factor $\exp(t/\pi_{\min})$) can be repeated with $\hat{\pi}$ instead. This yields $|R(\eta)|\leq C(\hat{\pi})\eta^{4}$ on $B_{\hat{\pi}}^{c}$. On $B_{\hat{\pi}}$, we just use the original bound $|R(\eta)|\leq C(\pi_{\min})\eta^{4}$. Plugging into the conditional expectation decomposition yields the desired result. ∎
Lemma˜C.3 gives us a way to shrink the remainder term ${\mathbb{E}}[R(\eta)]$ if we can find an appropriate $\mathcal{S}$. In the unclipped setting, we cannot benefit from this since it is hard to find $\mathcal{S}$}. However, this becomes useful in the upper-clipped case. We now proceed to the proof of Theorem˜4.3.
Theorem 4.3 (restated). Define $C_{i}:=\{A_{i}>0,\,r_{u}(\mathbf{y}^{(i)})>1+\varepsilon\}$. Let $\rho:=\mathbb{P}(C_{1})$ and $\delta={\mathbb{E}}[r_{u}(\mathbf{y}^{(1)})-(1+\varepsilon)\mid C_{1}]$. Then, for $\eta>0$ satisfying Eq.˜17 and any $p\in(\pi_{\min},1)$, we have
${\mathbb{E}}[{\mathcal{H}}(\pi_{\textnormal{new}}^{c})-{\mathcal{H}}(\pi_{\textnormal{old}})]\leq-c_{G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R_{c}(\eta)]+c(p)G\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right),$
where $c_{G}$ and $\Phi$ are defined as the same as in Theorem˜4.3, and
@@
We define ${\mathcal{S}}:=\{\mathbf{y}^{(i)}:A_{i}>0,\,r_{u}(\mathbf{y}^{(i)})\leq 1+\varepsilon\}$. Then, the bad event set can be set as follows,
$B_{\hat{\pi}}:=\{\exists\,i:A_{i}>0,\,r_{u}(\mathbf{y}^{(i)})\leq 1+\varepsilon,\,\pi_{\textnormal{old}}(\mathbf{y}^{(i)})\leq\hat{\pi}\}$.
By using Theorem˜4.1 and Lemma˜C.3, we have
${\mathbb{E}}[\Delta{\mathcal{H}}^{u}_{\textnormal{new}}]=-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R_{c}(\eta)].$ (18)
@@
By Lemma˜C.2, we have $\pi_{\textnormal{new}}^{c}(a)\leq(1+\varepsilon)\pi_{\textnormal{old}}(a)$ for every $a$. If $r_{u}(a)>1+\varepsilon$, we have
@@
$\log(\pi_{\textnormal{new}}^{u}(\mathbf{y}^{(i)}))=\log(\pi_{\textnormal{old}}(\mathbf{y}^{(i)}))r_{u}(\mathbf{y}^{(i)})\leq\min\{0,\log(pe^{\eta/(2\pi_{\min})})\}.$
Using Eq.˜19, restricting to indices $i$ where $C_{i}$ occurs, we have
@@
Conditioned on $\mathbf{y}^{(i)}=a$ and $\pi_{\textnormal{old}}(a)\leq p$, we have $K:=n_{\mathbf{y}^{(i)}}-1\sim\text{Binomial}(G-1,\pi_{\rm old}(a))$. Thus, we have
@@
where $X_{i}\leq X_{\max}:=\exp(\eta/(2\pi_{\min}))-(1+{\epsilon})$. This $X_{\max}$ is derived by recalling that $|\tilde{A}(a)|\leq 1/(2\pi_{\min})$ and noticing that Jensen’s inequality implies $Z(\eta):={\mathbb{E}}_{\pi_{\textnormal{old}}}[e^{\eta\tilde{A}(a)}]\geq e^{\eta{\mathbb{E}}_{\pi_{\textnormal{old}}}[\tilde{A}(a)]}=1$. Taking expectation of both sides of Eq.˜19 yields

### v2

# Exploration vs Exploitation: Rethinking RLVR through Clipping, Entropy, and Spurious Reward

@@ 4.1 One-step Policy Entropy Change under Random Rewards
${\mathbb{E}}[{\mathcal{H}}(\pi_{\textnormal{new}})-{\mathcal{H}}(\pi_{\textnormal{old}})]\leq-c_{G}\Phi(\pi_{\rm old})\eta^{2}+{\mathbb{E}}[R(\eta)]+c(p)G\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right),$
where $c_{G}$, $\Phi$ and $R(\eta)$ are defined as the same as in Theorem˜4.1; for other constants, see Section˜C.3.
@@ C.3 Missing Proofs in §4
Lemma C.1.
Let $\mathcal{V}$ be the vocabulary and $\tilde{A}$ be defined in Eq.˜3 with $L=1$. Then for $\eta>0$,
${\mathbb{E}}[{\mathcal{H}}(\pi_{\eta})-{\mathcal{H}}(\pi_{\rm old})]=-\tfrac{\eta^{2}}{2}{\mathbb{E}}[\mathrm{Var}_{\pi_{\rm old}}(\tilde{A})+\mathrm{Cov}_{\pi_{\rm old}}(\tilde{A}^{2},\log\pi_{\rm old})]+{\mathbb{E}}[R(\eta)],$ (17)
Whenever there exists $\hat{\pi}\in(0,1]$ such that $|\tilde{A}(a)|\leq\frac{1}{2\hat{\pi}}$ for all $a\in{\mathcal{V}}$, the remainder satisfies the explicit bound
$|{\mathbb{E}}[R(\eta)]|\leq\frac{e^{\eta/(2\hat{\pi})}}{24}\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176\eta}{\hat{\pi}}\right)\eta^{4}{\mathbb{E}}\!\left[{\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}^{4}]\right].$ (18)
Proof.
Write $L(a):=\log\pi_{\rm old}(a)$ and $\psi(\eta):=\log Z(\eta)$. Then $\log\pi_{\eta}(a)=L(a)+t\tilde{A}(a)-\psi(\eta)$ and hence
${\mathcal{H}}(\pi_{\eta})=-{\mathbb{E}}_{\pi_{\eta}}[\log\pi_{\eta}]=-{\mathbb{E}}_{\pi_{\eta}}[L]-\eta{\mathbb{E}}_{\pi_{\eta}}[\tilde{A}]+\psi(\eta).$ (19)
Differentiate $\pi_{\eta}(a)=\pi_{\rm old}(a)e^{\eta\tilde{A}(a)}/Z(\eta)$:
$\frac{\mathrm{d}}{\mathrm{d}\eta}\pi_{\eta}(a)=\pi_{\eta}(a)(\tilde{A}(a)-\psi^{\prime}(\eta)),\quad\psi^{\prime}(\eta)={\mathbb{E}}_{\pi_{\eta}}[\tilde{A}$].
Thus, for any $g:{\mathcal{V}}\to\mathbb{R}$,
[…]
From Eq.˜19 and $\psi^{\prime}(\eta)={\mathbb{E}}_{\pi_{\eta}}[\tilde{A}$], we can rewrite
${\mathcal{H}}(\pi_{\eta})=-{\mathbb{E}}_{\pi_{\eta}}[L]-\eta\psi^{\prime}(\eta)+\psi(\eta).$
[…]
Differentiate again and evaluate at $\eta=0$:
[…]
Now expand $\mathrm{Cov}_{\pi_{\eta}}(L,\tilde{A})={\mathbb{E}}_{\pi_{\eta}}[L\tilde{A}]-{\mathbb{E}}_{\pi_{\eta}}[L]{\mathbb{E}}_{\pi_{\eta}}[\tilde{A}$] and differentiate each expectation using Eq.˜20. A direct calculation gives
[…]
At $\eta=0$, since ${\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}]=0$, we have
$\left.\left(\frac{\mathrm{d}}{\mathrm{d}\eta}\mathrm{Cov}_{\pi_{\eta}}(L,\tilde{A})\right)\right|_{\eta=0}=\mathrm{Cov}_{\pi_{\rm old}}(L,\tilde{A}^{2}).$ (21)
Therefore
${\mathcal{H}}^{\prime\prime}(0)=-\mathrm{Var}_{\pi_{\rm old}}(\tilde{A})-\mathrm{Cov}_{\pi_{\rm old}}(\tilde{A}^{2},L),$ (22)
which, with expectation, is the coefficient of $\eta^{2}$ in Eq.˜17.
[…]
$f(\eta)=\frac{f^{\prime\prime}(0)}{2}\eta^{2}+\frac{f^{(4)}(\xi)}{24}\eta^{4}.$
Combining with Eq.˜22 gives Eq.˜17 with ${\mathbb{E}}[R(\eta)]=\frac{f^{(4)}(\xi)}{24}\eta^{4}$ and the bound
$\big|{\mathbb{E}}[R(\eta)]\big|\leq\frac{\eta^{4}}{24}\sup_{t\in[0,\eta]}|f^{(4)}(\eta)|\leq\frac{\eta^{4}}{24}\sup_{t\in[0,\eta]}{\mathbb{E}}\big[|{\mathcal{H}}^{(4)}(\eta)|\big].$ (23)
Thus, it remains to bound $|{\mathcal{H}}^{(4)}(\eta)|$ by ${\mathbb{E}}_{\pi_{\eta}}[\tilde{A}^{4}]$. Let $m_{\eta}:={\mathbb{E}}_{\pi_{\eta}}[\tilde{A}$], $X_{\eta}:=\tilde{A}-m_{\eta}$, and $L_{\eta}:=L-{\mathbb{E}}_{\pi_{\eta}}[L]$. A direct four-times differentiation of Eq.˜19 using Eq.˜20 yields the explicit identity
${\mathcal{H}}^{(4)}(\eta)$ $=-{\mathbb{E}}_{\pi_{\eta}}[L_{\eta}X_{\eta}^{4}]+4{\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{3}]{\mathbb{E}}_{\pi_{\eta}}[L_{\eta}X_{\eta}]+6{\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{2}]{\mathbb{E}}_{\pi_{\eta}}[L_{\eta}X_{\eta}^{2}]$
$\quad-3{\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{4}]+9\big({\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{2}]\big)^{2}+\eta\left(10\,{\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{2}]{\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{3}]-{\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{5}]\right).$ (24)
Now $\|L_{\eta}\|_{\infty}\leq\log(1/\pi_{\min})$ because $L(a)\in[\log\pi_{\min},0]$. Also, by Cauchy-Schwarz and Hölder,
$({\mathbb{E}}[X_{\eta}^{2}])^{2}\leq{\mathbb{E}}[X_{\eta}^{4}],\quad{\mathbb{E}}[|X_{\eta}|^{3}]{\mathbb{E}}[|X_{\eta}|]\leq{\mathbb{E}}[X_{\eta}^{4}].$
[…]
${\mathbb{E}}[|X_{\eta}|^{3}]\leq\frac{1}{\hat{\pi}}{\mathbb{E}}[X_{\eta}^{2}]\leq\frac{1}{\hat{\pi}}\sqrt{{\mathbb{E}}[X_{\eta}^{4}]},\quad{\mathbb{E}}[|X_{\eta}|^{5}]\leq\frac{1}{\hat{\pi}}{\mathbb{E}}[X_{\eta}^{4}],$
which gives
$\left|10{\mathbb{E}}[X_{\eta}^{2}]{\mathbb{E}}[X_{\eta}^{3}]-{\mathbb{E}}[X_{\eta}^{5}]\right|\leq 10{\mathbb{E}}[X_{\eta}^{2}]{\mathbb{E}}[|X_{\eta}|^{3}]+{\mathbb{E}}[|X_{\eta}|^{5}]\leq\frac{11}{\hat{\pi}}{\mathbb{E}}[X_{\eta}^{4}].$
Plugging these bounds into Section˜C.3 yields
$|{\mathcal{H}}^{(4)}(\eta)|\leq\left(12+11\log(\tfrac{1}{\pi_{\min}})+\tfrac{11t}{\hat{\pi}}\right){\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{4}].$ (25)
Finally, $(u-v)^{4}\leq 8(u^{4}+v^{4})$ and Jensen’s inequality imply ${\mathbb{E}}_{\pi_{\eta}}[X_{\eta}^{4}]\leq 16{\mathbb{E}}_{\pi_{\eta}}[\tilde{A}^{4}]$, hence
$|{\mathcal{H}}^{(4)}(\eta)|\leq\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176t}{\hat{\pi}}\right){\mathbb{E}}_{\pi_{\eta}}[\tilde{A}^{4}].$ (26)
By Jensen’s inequality and the fact that ${\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}]=0$, we have $Z(\eta)\geq 1$. Hence for $\eta\geq 0$,
${\mathbb{E}}_{\pi_{\eta}}[\tilde{A}^{4}]=\frac{1}{Z(\eta)}{\mathbb{E}}_{\pi_{\rm old}}[e^{t\tilde{A}}\tilde{A}^{4}]\leq{\mathbb{E}}_{\pi_{\rm old}}[e^{t\tilde{A}}\tilde{A}^{4}]\leq e^{t/(2\hat{\pi})}{\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}^{4}],$
Combining with Eq.˜26 gives, for $t\in[0,\eta]$,
${\mathbb{E}}[|{\mathcal{H}}^{(4)}(\eta)|]\leq e^{t/(2\hat{\pi})}\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176t}{\hat{\pi}}\right){\mathbb{E}}\!\left[{\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}^{4}]\right].$
Taking the supremum over $t\in[0,\eta]$ yields
$\sup_{t\in[0,\eta]}{\mathbb{E}}[|{\mathcal{H}}^{(4)}(\eta)|]\leq e^{\eta/(2\hat{\pi})}\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176\eta}{\hat{\pi}}\right){\mathbb{E}}\!\left[{\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}^{4}]\right].$
Plug this into Eq.˜23 to obtain Eq.˜18. ∎
Corollary C.2.
Under the same setting as in Lemma˜C.1, if we know that conditioned on an event $B$, $|\tilde{A}(a)|\leq\frac{1}{2\hat{\pi}}$, then we have a refined bound
$|{\mathbb{E}}[R(\eta)]|\leq\mathbb{P}(B)C(\hat{\pi})+(1-\mathbb{P}(B))C(\pi_{\min}),$ (27)
@@
$C(\pi):=\frac{e^{\eta/(2\pi)}}{24}\left(192+176\log(\tfrac{1}{\pi_{\min}})+\tfrac{176\eta}{\pi}\right){\mathbb{E}}\!\left[{\mathbb{E}}_{\pi_{\rm old}}[\tilde{A}^{4}]\right]\eta^{4}.$
@@
Let $q:=\mathbb{P}(B)$. By the tower property,
${\mathbb{E}}[R(\eta)]=q{\mathbb{E}}[R(\eta)\mid B]+(1-q){\mathbb{E}}[R(\eta)\mid B^{c}],$
hence by the triangle inequality,
$|{\mathbb{E}}[R(\eta)]|\leq q|{\mathbb{E}}[R(\eta)\mid B]|+(1-q)|{\mathbb{E}}[R(\eta)\mid B^{c}]|.$
@@
Lemma C.3.
Let $\mathcal{V}$ be the vocabulary and $\tilde{A}$ be defined in Eq.˜3 with $L=1$, then
${\mathbb{E}}[\textnormal{Var}_{\pi_{\textnormal{old}}}(\tilde{A})]=\tfrac{1-2^{1-G}}{G}(|{\mathcal{V}}|-1),$
${\mathbb{E}}[\textnormal{Cov}_{\pi_{\textnormal{old}}}(\tilde{A}^{2},\log(\pi_{\textnormal{old}}))]=\tfrac{1-2^{1-G}}{G}\left(\sum_{a}L(a)-|{\mathcal{V}}|\mathbb{E}_{\pi_{\rm old}}[L])\right)$
where $S_{1},S_{2}>0$ are
$S_{1}=\tfrac{|{\mathcal{V}}|-1}{\pi_{\min}}+\tfrac{1}{1-(|{\mathcal{V}}|-1)\pi_{\min}},\quad S_{2}=\tfrac{|{\mathcal{V}}|-1}{(\pi_{\min})^{2}}+\tfrac{1}{(1-(|{\mathcal{V}}|-1)\pi_{\min})^{2}}.$
Proof.
@@
Finally, we compute ${\mathbb{E}}[{\mathbb{E}}_{\pi_{\textnormal{old}}}[\tilde{A}^{4}]]$. We define $I_{i}(a)=\bm{1}\{\mathbf{y}^{(i)}=a\}$. Thus, we have $I_{i}(a)\sim\textnormal{Bernoulli}(p)$ i.i.d. and it is independent of $\{A_{i}\}_{i=1}^{G}$. We write $S(a)=\sum_{i=1}^{G}A_{i}I_{i}(a)$ and
@@
$S_{1}=\tfrac{|{\mathcal{V}}|-1}{\pi_{\min}}+\tfrac{1}{1-(|{\mathcal{V}}|-1)\pi_{\min}},\quad S_{2}=\tfrac{|{\mathcal{V}}|-1}{(\pi_{\min})^{2}}+\tfrac{1}{(1-(|{\mathcal{V}}|-1)\pi_{\min})^{2}}.$
${\mathbb{E}}[{\mathcal{H}}(\pi_{\textnormal{new}})-{\mathcal{H}}(\pi_{\textnormal{old}})]=-c_{G}\Phi(\pi_{\textnormal{old}})\eta^{2}+\mathbb{E}[R(\eta)],\quad|R(\eta)|\leq C\eta^{4}.$
where $c_{G}:=\tfrac{1-2^{1-G}}{2G}$ and
with
Proof.
We rewrite the update rule as follows,
We define $\psi(\eta):=\log Z(\eta)$ and $u(a):=\eta\tilde{A}(a)-\psi(\eta)$. In Lemma˜2.2, we have shown $|\tilde{A}(a)|\leq 1/(2\pi_{\min})$ for all $a$, so by taking $\hat{\pi}:=\pi_{\min}$, we can apply Lemma˜C.1 to obtain
${\mathbb{E}}[{\mathcal{H}}(\pi_{\eta})-{\mathcal{H}}(\pi_{\rm old})]=-\tfrac{\eta^{2}}{2}{\mathbb{E}}[\mathrm{Var}_{\pi_{\rm old}}(\tilde{A})+\mathrm{Cov}_{\pi_{\rm old}}(\tilde{A}^{2},\log\pi_{\rm old})]+{\mathbb{E}}[R(\eta)].$
Finally, we apply Lemma˜C.3 and define
to obtain the desired result. ∎
@@
Finally, we show that $r_{c}(a)\leq 1+\varepsilon$ for all $a\in{\mathcal{V}}$ given Eq.˜28. Suppose that there are some $a^{\star}$ so that $r_{c}(a^{\star})>1+\varepsilon$. Then, by Lemma˜C.4, we have $\xi_{a^{\star}}=0$ and
@@
Putting these pieces together yields a direct contradiction to Eq.˜28 as follows,
@@
Theorem 4.3 (restated). Define $C_{i}:=\{A_{i}>0,\,r_{u}(\mathbf{y}^{(i)})>1+\varepsilon\}$. Let $\rho:=\mathbb{P}(C_{1})$ and $\delta={\mathbb{E}}[r_{u}(\mathbf{y}^{(1)})-(1+\varepsilon)\mid C_{1}]$. Then, for $\eta>0$ satisfying Eq.˜28 and any $p\in(\pi_{\min},1)$, we have
${\mathbb{E}}[{\mathcal{H}}(\pi_{\textnormal{new}}^{c})-{\mathcal{H}}(\pi_{\textnormal{old}})]\leq-c_{G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R(\eta)]+c(p)G\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right),$
where $c_{G}$, $\Phi$ and $R(\eta)$ are defined as the same as in Theorem˜4.1, and
@@
By first part of Lemma˜C.1 and Lemma˜C.3, we have
${\mathbb{E}}[\Delta{\mathcal{H}}^{u}]=-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R(\eta)].$ (29)
@@
By Lemma˜C.5, we have $\pi_{\textnormal{new}}^{c}(a)\leq(1+\varepsilon)\pi_{\textnormal{old}}(a)$ for every $a$. If $r_{u}(a)>1+\varepsilon$, we have
@@
$\log(\pi_{\textnormal{new}}^{u}(\mathbf{y}^{(i)}))=\log(\pi_{\textnormal{old}}(\mathbf{y}^{(i)})r_{u}(\mathbf{y}^{(i)}))\leq\min\{0,\log(pe^{\eta/(2\pi_{\min})})\}.$
Using Eq.˜30, restricting to indices $i$ where $C_{i}$ occurs, we have
@@
Define $K:=n_{\mathbf{y}^{(i)}}-1$. Conditioned on $\mathbf{y}^{(i)}=a$ and $\pi_{\textnormal{old}}(a)\leq p$, we have
$K\leq\sum_{j=1}^{G}\bm{1}\{\mathbf{y}^{(j)}=a\}-1\sim\text{Binomial}(G-1,\pi_{\rm old}(a)).$
Thus, we have
@@
where $X_{i}\leq X_{\max}:=\exp(\eta/(2\pi_{\min}))-(1+{\epsilon})$. This $X_{\max}$ is derived by recalling that $|\tilde{A}(a)|\leq 1/(2\pi_{\min})$ and noticing that Jensen’s inequality implies $Z(\eta):={\mathbb{E}}_{\pi_{\textnormal{old}}}[e^{\eta\tilde{A}(a)}]\geq e^{\eta{\mathbb{E}}_{\pi_{\textnormal{old}}}[\tilde{A}(a)]}=1$. Taking expectation of both sides of Eq.˜30 yields